An insurance company crashed four cars of the same model at 5 miles per hour. The costs of repair for each of the four crashes were ​$ 409​, ​$428, ​$457​, and ​$237 . Compute the​ mean, median, and mode cost of repair
Answer #1

Solution: Let the first write the given things. The cost of repair for each of the four crashes were $409, $428,$457,and $237.

Here xi=409,428,457 and 237 the values of repair cost

n=4 .number of crashes cars.

First thing calculate is the Mean.

Mean= \frac{\sum_{i=1}^{n}xi}{n} = (409+428+457+237)/4 = 382.75

Second thing is Median

For median we have to set the cost of repair in ascending order.

The order is 237 , 409, 428, 457

Here we have given four values. If the observation are even the we add middle most values and divide by 2

Median = (409+428)/2 = 418.5

The last thing is Mode.

Here Mode does not exist . Because of no cost repeated two times. Mode only exist when any value repeated frequently.